# OmniGenome: Aligning RNA Sequences with Secondary Structures in Genomic Foundation Models

### v2

@@ 4 Related Works
Biological sequence modelling, including DNA, RNA, and protein, has attracted attention in recent years. Protein modelling, e.g., AlphaFold [36, 37, 38] and ESM [39], has been studied for many years compared to DNA and RNA modelling. In the realm of genomic sequence modelling, several early works aimed at addressing diversified genome downstream subtasks. For instance, DNABERT [40] adapts the architecture of BERT [27] for genomic sequence modelling, showing preliminary performance for in-silico genomic tasks. DNABERT2 [22], a multi-species FM improved based on DNABERT, proposes replacing k-mers tokenization with BPE tokenization to improve model performance. To explore the performance of large-scale FMs, the nucleotide transformers (V1 & V2) [26], AgroNT [20] and SegmentNT [41] leveraged billions of parameters to boost genomic sequence modelling and achieved promising performance in understanding DNA genome, with model scales up to billion and billion parameters, respectively. Agro-NT [20] was pre-trained on multi-species edible plant DNA sequences but failed to transfer effectively to RNA sequence modelling in our experiments. To address the modelling capacity problem caused by the remarkable lengths of genomes, there is a growing focus on the necessity of long-range sequence modelling and the introduction of autoregressive FMs, namely, HyenaDNA [1] and Evo [42]. In terms of RNA genomic modelling, there have been some preliminary works, such as scBERT [43], RNABERT [44], RNA-FM [12], RNA-MSM [45], and RNAErnie [46], to name a few. However, these methods have only trained the FMs on a limited-scale database, as RNA sequences are generally expensive to obtain. Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) [47], 5’ untranslated regions (5’UTR) [35], 3’ untranslated regions (3’UTR) [25], or precursor mRNA sequences [24], thus limiting the models’ ability to capture the diversity of RNA sequences. Uni-RNA [34] has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM [35] adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training. However, these FMs are not available for Str2Seq mapping and suffer from limited model and data scales that fail to uncover the comprehensive efficacy of sequence-structure alignment on a wide set of genomic tasks. ERNIE-RNA [48] feeds the RNA structure along with the sequence into the model and improves the downstream tasks. However, it also ignores the significance of Str2Seq prediction capability. In a nutshell, existing FMs fail to achieve sequence-structure alignment without exception.

### v3

@@ 4 Related Works
Current RNA FMs focused on sequence-to-structure mapping, e.g., end-to-end secondary structure prediction. However, to the best of our knowledge, the sequence-structure alignment in RNA genome modelling has yet been investigated in the literature. There have been some preliminary works, such as scBERT (Yang et al. 2022), RNABERT (Akiyama and Sakakibara 2022), RNA-FM (Chen et al. 2022), RNA-MSM (Zhang et al. 2023), and RNAErnie (Wang et al. 2024), to name a few. However, these methods have only trained the FMs on a limited-scale database, as RNA sequences are generally expensive to obtain. Some FMs focus on specific types of RNA sequences, such as coding sequences (CDS) (Hallee, Rafailidis, and Gleghorn 2023), 5’ untranslated regions (5’UTR) (Chu et al. 2024), 3’ untranslated regions (3’UTR) (Yang et al. 2023), or precursor mRNA sequences (Chen et al. 2023), thus limiting the models’ ability to capture the diversity of RNA sequences. Uni-RNA (Wang et al. 2023) has been reported to achieve good performance due to the large scale of the model and database, however, it is not open-sourced and cannot be compared in the experiments.
In short, the existing RNA FMs neglect the significance of sequence-structure alignment in RNA genome modelling, while the 5UTR-LM (Chu et al. 2024) adopts the secondary structure prediction as a pre-training objective to achieve Seq2Str prediction in pre-training. However, these FMs are not available for Str2Seq mapping and suffer from limited model and data scales that fail to uncover the comprehensive efficacy of sequence-structure alignment on a wide set of genomic tasks. ERNIE-RNA (Yin et al. 2024) feeds the RNA structure along with the sequence into the model and improves the downstream tasks. However, it also ignores the significance of Str2Seq prediction capability. In a nutshell, existing FMs fail to achieve sequence-structure alignment without exception.